Suppose that the random variable X represents the length of a punched part in meters. Let Y be the length of the part in inches. Please report your answer in 1 decimal place.

  • If V(X) = 0.29, what is the variance of Y?

1 m = 39.37 inches
Answer #1

Solution:

We have

1 m = 39.37 inches

That is, we have

Y = 39.37*X

We know that

If Y = a*X, then V(Y) = a^2*Var(X)

So, we have

V(Y) = 39.37^2*V(X)

We are given V(X) = 0.29

V(Y) = 39.37^2*0.29

V(Y) = 449.4991

V(Y) = 449.5

Answer: Variance of Y = 449.5